Consider an economy with two firms. Firm A produces orange and Firm B produces orange juice. In a given year, Firm A produces 10 million pounds oranges,ssio pounds of these oranges to Firm B. sells 3 million pounds oranges to domestic consumers, and exports 3 ш1llion pounds oranges. Each pound of orange is sold at S1. Firm A pays S6 million in wages to domestic workers and S1 on tax to the government. Firm B produces 4 million boxes of orange juice. In addition, they import 2 million boxes of orange juice. The import price of orange juice is $2 per box. 5 million boxes of orange juice are sold to domestic consumers at $3 per box, and the remaining 1 million boxes of orange juice is stored in the inventory. Firm B pays S5 million iın wages to the domestic workers, and pays S2 million in taxes to the government. The government provides education for the domestic consumers, and the cost of the educational services is just the wages paid to the teachers, which is S4 lon. Finish the following tasks and briefly show the steps. 1.1 Calculate the GDP using product approach. 1.2 Calculate the GDP using expenditure approach. 1.3 Calculate the GDP using income approach.
Answer #1

Product Method.

Total Production of oranges by Firm A =10 million pounds

Price per pound =$1

Value of the total product=10 x 1 =$10 million.

Tax=$1 million

Value added=10 million- 1 =$9 million

Firm B Value of Production of juice=4 million x$3=$12 million

Tax=$2 Million.

Net value added=$12 -2 =10 million

Total value added Firm A+ B= 10 million +9 million=$ 19 million

Income Method

Firm A. Wage=$6 million

Firm B.Wage=$ 5 million

Income of teachers=$ 4 million

Export Price=3 x 1=$3 million

Import price=2 x 2=$4 m.

Net factor income=3-4= -1 m.dollars

Net Income=15 - 1=$14 m.

Expenditure method

sale to Firm B by A= 4 x 1=$4 m.

Domestic sale of A= 3 x 1=$3 m

Sale of juice by B=5 x 3= $15 m.-$4( intermediary)=11 m.dollars

Government expenditure=$ 4 m.

Net factor income=Export - Import=3-4=( -1)

Tax=$1 m.+ $2.m= $3 m.

Net Income=(4+3 +11+4) -(3-1)=20